OpenStudy (anonymous):

2(h-8)-h=h-16 h=8 h=-8 infinitely many solutions no solution

OpenStudy (azureilai):

2h-16-h=h-16 2h-16-h-h=-16 -16=-16 h can be anything

OpenStudy (agent0smith):

First distribute: 2(h-8)-h=h-16 becomes 2h-16-h=h-16 now see if you can combine like terms.